how to find the amplitude of y = 0.8/[(4x+5t)^2+5]
1
Expert's answer
2020-05-07T09:50:14-0400

As per the given question,

"y =\dfrac{ 0.8}{[(4x+5t)^2+5]}" here y is the displacement about the equilibrium,

So, displacement will be maximum, if denominator will be minimum,

"\Rightarrow [(4x+5t)^2]=0"

"\Rightarrow 4x+5t=0"

"x=-5t\/4"

"y=\dfrac{0.8}{5}=0.16"



at t=0,